Recording Inventory Costs [LO 4, 5] Vulcan Molding produces molded rubber components. At the start of the year, the company estimated that it would incur $3,500,000 of direct labor cost and $9,000,000 of manufacturing overhead. Overhead is allocated to production on the basis of direct labor cost. Actual materials used during the year were $5,750,000, actual direct labor cost was $4,000,000, and actual overhead was $11,000,000.

REQUIRED

  1. a. Calculate the overhead rate for the current year. Round to 2 decimal places.
  2. b. Prepare the journal entry to record use of direct material.
  3. c. Prepare the journal entry to record direct labor.
  4. d. Prepare the journal entry to record manufacturing overhead applied to production.
  5. e. Prepare the journal entry to close the balance in manufacturing overhead to cost of goods sold.
Answer #1

Please provide rating..

Answer a)
Calculate the overhead rate for the current year
estimated direct labor $3,500,000
Estimated overhead $9,000,000
Overhead rate = 9000000/3500000 $2.57 per direct labor cost
Dr Cr
answer b) work in process $575,000
        Direct materials $575,000
answer c) Work in process $4,000,000
        Direct labor $4,000,000
answer d) work in process =4000000*2.57 $10,285,714
      Manufacturing overheads $10,285,714
answer e) cost of goods sold 714286
      Manufacturing oveheads $714,286
=11000000-10285714